# ELASTIC: Efficient Once For All Iterative Search for Object Detection on Microcontrollers

### v2

@@ Abstract
Deploying high-performance object detectors on TinyML platforms poses significant challenges due to tight hardware constraints and the modular complexity of modern detection pipelines. Neural Architecture Search (NAS) offers a path toward automation, but existing methods either restrict optimization to individual modules—sacrificing cross-module synergy—or require global searches that are computationally intractable. We propose ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a unified, hardware-aware NAS framework that alternates optimization across modules (e.g., backbone, neck, and head) in a cyclic fashion. ELASTIC introduces a novel Population Passthrough mechanism in evolutionary search that retains high-quality candidates between search stages, yielding faster convergence, up to an 8% final mAP gain, and eliminates search instability observed without population passthrough. In a controlled comparison, empirical results show ELASTIC achieves +4.75% higher mAP and 2× faster convergence than progressive NAS strategies on SVHN, and delivers a +9.09% mAP improvement on PascalVOC given the same search budget. ELASTIC achieves 72.3% mAP on PascalVOC, outperforming MCUNET by 20.9% and TinyissimoYOLO by 16.3%. When deployed on MAX78000/MAX78002 microcontrollers, ELASTIC-derived models outperform Analog Devices’ TinySSD baselines, reducing energy by up to 71.6 %, lowering latency by up to 2.4, and improving mAP by up to 6.99 percentage points across multiple datasets.
@@ I Introduction
Object detection is a fundamental task in computer vision with real-world applications in mobile robotics, autonomous vehicles, smart surveillance, and embedded systems [1, 2]. While modern detectors achieve remarkable accuracy and speed on powerful hardware, their deployment on TinyML platforms—microcontrollers and edge devices with sub-MB memory, ultra-low power, and limited compute budgets—remains challenging [3, 4, 5, 6, 7].
However, unlike image classification, object detection involves multi-stage pipelines—typically composed of a backbone [8], neck [9], and head [1, 2]—each contributing uniquely to accuracy, latency, and memory footprint [10]. These interdependent components enlarge the design space exponentially, complicating deployment on microcontrollers, where model size, SRAM usage, and inference cycles must be tightly controlled [11, 5, 12].
Neural Architecture Search (NAS) offers a compelling approach to automate architecture design under tight resource constraints [13, 14]. Constraint-aware NAS has advanced edge classification [15, 3, 16], but object detection adds new layers of complexity. Prior detection NAS methods focus on individual modules (e.g., backbones [17], necks [9], or heads [18]), often ignoring cross-module dependencies. Full-pipeline NAS is possible [19], but scales poorly due to the massive combinatorial space—e.g., over configurations in realistic search settings [18, 20].
@@
Unified Iterative NAS: We propose a novel NAS framework that cyclically optimizes different modules (i.e. backbone, neck, head), capturing inter-module interactions and expanding the search space. Experimental results demonstrate that our approach achieves architectures with up to 4.75% higher mAP than progressive search strategies in early stages while being twice as fast.
Population Passthrough Mechanism: We develop a memory-based population retention scheme that mitigates performance drops during module alternation. On PascalVOC, this mechanism improves final mAP from 22.1% without passthrough to 30.83% and stabilizes search convergence across iterations.
Enhanced Search Space: ELASTIC significantly enhances the search space over time. On SVHN, the mean mAP of randomly sampled architectures improves from 44.61% (joint) to 70.68% after 5 iterations, while variance reduces by 89.7%. On PascalVOC, the mean mAP improves from 4.99% to 28.32%, with a 95.3% variance reduction.
Once-for-All Deployment on Resource-Constrained Devices: We validate ELASTIC in designing versatile models tailored to various resource constraints, demonstrating their inference efficiency on the MAX78000 platform. Our ELASTIC model significantly outperforms the baseline [21] on the MAX78000 platform, achieving a 45.4% reduction in energy consumption, a 29.3% decrease in latency, and a 4.5% improvement in mAP, demonstrating enhanced computational efficiency and accuracy.
@@ II Related Work
NAS for Object Detection. Object detection architectures are modular, comprising a backbone, neck, and head [2, 1, 10]. NAS methods have targeted these components individually or jointly: DetNAS [17] searches backbone architectures via a one-shot supernet; NAS-FPN [9] focuses on FPN design; NAS-FCOS [18] extends search to the FPN and head using reinforcement learning; MobileDets [22] introduces latency-aware backbone optimization for mobile accelerators; and Hit-Detector [19] jointly searches all modules but incurs high computational cost. While effective on larger systems, these methods are not designed for deployment on resource-constrained microcontrollers.
@@ III Methodology
We propose ELASTIC, a novel constrained NAS strategy that iteratively optimizes individual modules while maintaining inter-module context. This approach balances search efficiency with architectural synergy, enabling the discovery of compact, high-performing object detectors for real-time deployment on microcontrollers. To further improve convergence, we introduce a Population Passthrough mechanism that preserves top-performing architecture candidates across iterations during evolutionary search.
@@ III-B Iterative Evolutionary Architecture Search
At the core of our ELASTIC framework is an Iterative Evolutionary Architecture Search, which alternates optimization between modular components—specifically the backbone and the head—over successive cycles. Unlike conventional evolutionary NAS that operates in a single global pass, our method applies modular evolutionary search iteratively, enabling progressive refinement and co-adaptation across modules. This design explicitly captures cross-module interactions while maintaining search tractability under TinyML constraints.
The evolutionary procedure within each iteration follows these steps:
Population Initialization: The candidate population is composed of two parts: (1) a set of high-performing architectures retained from the module’s previous iteration via Population Passthrough, and (2) newly sampled architectures that maintain diversity in the search.
Genetic Variation. We generate new architectural candidates using crossover and mutation, subject to module-specific resource constraints (e.g., for the backbone and for the head). All offspring are filtered to ensure constraints are satisfied.
Fitness Evaluation: Each candidate is instantiated within the weight-shared supernet and evaluated using mean average precision (mAP) on a validation set. Since only one module is varied at a time, the evaluation reflects the candidate’s performance in the context of a fixed complementary module.
Selection: The highest-performing candidates are retained to seed the next generation. This ensures that architectural refinements are preserved and that future iterations start from strong initial configurations.
After optimizing the current module (e.g., backbone), the algorithm switches focus to the other module (e.g., head) and repeats the process. This iterative back-and-forth search continues for a predefined number of cycles or until convergence criteria are met. This structured alternation between modules allows ELASTIC to gradually co-adapt modular components under deployment constraints.
@@ III-C Population Passthrough
A central challenge in iterative modular NAS is maintaining optimization continuity across alternating search phases. When a search process shifts focus from one module (e.g., backbone) to another (e.g., head), conventional evolutionary algorithms often discard the entire population from the previous cycle. This reinitialization disrupts search momentum, leading to slower convergence and redundant exploration.
@@
| SVHN | Backbone-Only [17] | 75.53% | +0.18% | 10.8 hrs | 137.6 M | 0.73 M | 2.62 ms |
| Head-Only [18] | 75.35% | +0.00% | 4.7 hrs | 475.3 M | 1.01 M | 2.69 ms | |
@@
To mitigate this, we introduce Population Passthrough, a memory-based mechanism that explicitly retains high-quality search knowledge across module alternations. Each module maintains a passthrough memory buffer , which stores a ranked set of elite candidates based on performance from its most recent optimization cycle. When control returns to module in a later iteration, the initial population for evolutionary search is composed as follows:
Elite Inheritance: A partition of the population is sampled directly from , preserving high-performing architectural patterns discovered in earlier cycles.
Exploration: The remaining candidates are randomly sampled from the current search space, ensuring sufficient diversity to explore new regions and avoid premature convergence.
Notably, Population Passthrough differs from traditional aging [14], which typically operate within a single module or generation. Instead, it functions across alternating optimization cycles in a modular setting, making it uniquely suited for constrained multi-part architectures such as those required by TinyML.
@@ III-D Resource Allocation and Search Budget
In addition, ELASTIC is designed to support search interruption, allowing the iterative process to be paused at any stage and a valid, high-quality subnetwork to be extracted. This property is particularly beneficial in real-world TinyML development workflows, where compute resources are limited and iterative refinement may need to occur incrementally [4]. The ability to recover performant models mid-search adds flexibility and practicality to constrained NAS pipelines.
@@ IV-A Experimental Setup
We construct four OFA-based supernets tailored for diverse TinyML-relevant datasets: (i) SVHN, (ii) VGGFace2, (iii) a five-class subset of PascalVOC, and (iv) the full PascalVOC dataset. These datasets are selected to cover a range of TinyML application domains, from lightweight digit recognition to face identification and general-purpose object detection, while avoiding the prohibitive memory and compute demands of large-scale datasets such as COCO.
The SVHN, VGGFace2, and PascalVOC subset supernets adopt a one-stage SSD-style architecture, while the full PascalVOC supernet adopts a MobileNet–YOLO style design that combines a lightweight MobileNet backbone [8] with YOLO-style detection heads [1] to accommodate the increased dataset complexity. For the SSD-style supernets, the search space varies (i) layer widths, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of layers allocated at each feature map. For the MobileNet–YOLO supernet, the search space varies (i) expansion ratios in inverted bottlenecks, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of inverted bottleneck blocks per feature map.
@@
Evolutionary search is performed with a population size of 100, mutation probability of 0.2, mutation ratio of 0.5, and a parent selection ratio of 0.25. Each search is budgeted for 55 generations on SVHN and VGGFace2, and 30 generations on the PascalVOC subset and full PascalVOC datasets, respectively. Convergence is defined as the point at which the best candidate architecture’s mAP is within 1% of the final value achieved at the end of the search.
@@ IV-B Comparative Search Evaluation
To assess ELASTIC’s suitability for microcontrollers, we extend our experiments using the MCUNET [3] search space with the neck module included for fair comparison. Benchmarked on the full PascalVOC () dataset with similar parameter budgets (Table II), our ELASTIC-derived detector achieves 72.3% mAP, surpassing MCUNET by 20.9%, TinyissimoYOLO’s [5] best model by 16.3%, and MCUNetV2-M4/H7 [25] by 7.7% and 4.0%, respectively, while using 50–75% fewer MACs. Unlike MCUNetV2 [25], which requires external patch-based inference, ELASTIC supports native end-to-end inference, making it more general and easier to deploy across diverse microcontrollers.
@@
IV-C Impact of Population Passthrough on Search Stability and Efficiency
Iterative NAS often suffers from performance drops when switching between modules due to population reinitialization. Without memory, each alternation incurs a “catch-up” phase, degrading search stability and slowing convergence. To address this, we introduce Population Passthrough, a memory mechanism that retains top-performing architectures across module alternations.
Figure 3 shows the effect of varying passthrough ratios on the PascalVOC dataset. Without passthrough, mAP repeatedly drops after each module switch and saturates below 22%. In contrast, ratios of 50–80% yield smoother trajectories and reach over 30% mAP within half the iterations. The best results are obtained with a 60% passthrough ratio, which strikes an effective balance between inheritance of strong candidates and exploration of new architectures. Higher ratios risk overfitting to stale subnets, while lower ratios lead to instability.

### v3

@@ Abstract
Deploying high-performance object detectors on TinyML platforms poses significant challenges due to tight hardware constraints and the modular complexity of modern detection pipelines. Neural Architecture Search (NAS) offers a path toward automation, but existing methods either restrict optimization to individual modules—sacrificing cross-module synergy—or require global searches that are computationally intractable. We propose ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a unified, hardware-aware NAS framework that alternates optimization across modules (e.g., backbone, neck, and head) in a cyclic fashion. ELASTIC introduces a novel Population Passthrough mechanism in evolutionary search that retains high-quality candidates between search stages, yielding faster convergence, up to an 8% final mAP gain, and eliminates search instability observed without population passthrough. In a controlled comparison, empirical results show ELASTIC achieves +4.75% higher mAP and 2× faster convergence than progressive NAS strategies on SVHN, and delivers a +9.09% mAP improvement on PascalVOC given the same search budget. ELASTIC achieves 72.3% mAP on PascalVOC, outperforming MCUNET by 20.9% and TinyissimoYOLO by 16.3%. When deployed on MAX78000/MAX78002 microcontrollers, ELASTIC-derived models outperform Analog Devices’ TinySSD baselines, reducing energy by up to 71.6 %, lowering latency by up to 2.4, and improving mAP by up to 6.99 percentage points across multiple datasets. The experimental videos and codes are available on the project website111Project Website: https://nail-uh.github.io/elastic.github.io/.
@@ I Introduction
Object detection is a fundamental task in computer vision with real-world applications in mobile robotics, autonomous vehicles, smart surveillance, and embedded systems. While modern detectors achieve remarkable accuracy and speed on powerful hardware, their deployment on TinyML platforms—microcontrollers and edge devices with sub-MB memory, ultra-low power, and limited compute budgets—remains challenging [13, 11, 15, 17, 25].
Unlike image classification, object detection relies on a multi-stage pipeline—backbone, neck [7], and head—whose design choices are strongly interdependent: modifying one module changes the feature distributions and optimal configurations of the others. This cross-module coupling makes the detector design space highly combinatorial, which is particularly challenging on microcontrollers where flash/SRAM budgets and inference cycles must be tightly controlled [21, 15, 1]. Recent hardware-aware NAS efforts for pure MCU targets have largely focused on image classification [5, 6], where the optimization is over single-stream networks with a single prediction head and classification objectives; in contrast, MCU detection requires jointly reasoning over the coupled backbone–neck–head pipeline and mAP-driven evaluation.
Recent TinyML-oriented detection research improves efficiency via specialized architectures and deployment-aware design. Hardware-conscious model and compilation/runtime co-design enable practical MCU execution [13, 1, 21], while lightweight detectors and variants (including transformer-inspired designs) explore accuracy–efficiency trade-offs under tight energy/latency constraints [15, 17, 25, 11]. Despite these advances, selecting an optimal detector for a specific MCU remains non-trivial because accuracy, flash/SRAM footprint, and inference cycles are tightly coupled, and manual design or ad-hoc scaling rarely navigates these constraints systematically at the full detector level.
Neural Architecture Search (NAS) offers a compelling approach to automate architecture design under tight resource constraints. Constraint-aware NAS has advanced edge classification [2, 13, 14], but object detection adds new layers of complexity. Prior detection NAS methods focus on individual modules (e.g., backbones [4], necks [7], or heads [22]), often ignoring cross-module dependencies. Full-pipeline NAS is possible [8], but scales poorly due to the massive combinatorial space—e.g., over configurations in realistic search settings [22, 19]. These observations expose a key gap: MCU-targeted detection NAS must preserve cross-module co-adaptation (backbone–neck–head interactions) without incurring the prohibitive cost of full end-to-end search, while enforcing hard resource constraints so discovered models are directly deployable.
@@
Unified Iterative NAS: We propose a cyclic module-wise NAS framework that alternates optimization across the backbone, neck, and head to preserve cross-module co-adaptation while reducing search cost by restricting the active search dimension at each iteration. In our early-stage PascalVOC search, ELASTIC achieves up to +4.75% mAP over the progressive baseline while on SVHN, ELASTIC reduces search cost until convergence from 30.8 to 12.5 GPU hours.
Population Passthrough Mechanism: We introduce Population Passthrough, an elite-carryover strategy that stabilizes the iterative module alternation by retaining top-performing candidates when switching the active search module, rather than reinitializing the evolutionary population. This directly mitigates the abrupt performance drops observed under naive module switching and enables consistent convergence across iterations. On PascalVOC, enabling passthrough improves the final mAP from 22.1% (iterative search without passthrough) to 30.83% (with passthrough), demonstrating that the mechanism is a necessary component for making cyclic multi-module optimization effective in practice.
Enhanced Search Space: ELASTIC enhances the search space over time. On SVHN, the mean mAP of randomly sampled architectures improves from 44.61% (joint) to 70.68% after 5 iterations, while variance reduces by 89.7%. On PascalVOC, the mean mAP improves from 4.99% to 28.32%, with a 95.3% variance reduction.
Once-for-All Deployment on Resource-Constrained Devices: We validate ELASTIC in designing versatile models tailored to various resource constraints, demonstrating their inference efficiency on the MAX78000/02 and STM32F746 platform. Our ELASTIC model significantly outperforms the baseline [9] on the MAX78000 platform, achieving a 45.4% reduction in energy consumption, a 29.3% decrease in latency, and a 4.5% improvement in mAP, demonstrating enhanced computational efficiency and accuracy.
@@ II Related Work
NAS for Object Detection. Object detection architectures are modular, comprising a backbone, neck, and head. NAS methods have targeted these components individually or jointly: DetNAS [4] searches backbone architectures via a one-shot supernet; NAS-FPN [7] focuses on FPN design; NAS-FCOS [22] extends search to the FPN and head using reinforcement learning; MobileDets [23] introduces latency-aware backbone optimization for mobile accelerators; and Hit-Detector [8] jointly searches all modules but incurs high computational cost. While effective on larger systems, these methods are not designed for deployment on resource-constrained microcontrollers.
@@ III Methodology
This section introduces ELASTIC, a constrained iterative NAS method for TinyML object detection that explicitly treats the detector as a multi-module pipeline (backbone–neck–head). The key distinction from prior NAS strategies is that ELASTIC performs cyclic module-wise optimization: at each iteration, we optimize one module while keeping the others fixed, and we revisit modules multiple times to enable cross-module co-adaptation without incurring the prohibitive cost of full end-to-end search. However, naively switching the optimized module can destabilize the search and cause abrupt drops in validation performance. To address this, we propose Population Passthrough, which retains top-performing architecture candidates at the end of each module stage and injects them into the next stage, stabilizing convergence across iterations under explicit resource constraints.
@@ III-B Iterative Evolutionary Architecture Search
At the core of ELASTIC is an Iterative Evolutionary Architecture Search that alternates optimization between the backbone and the head over successive cycles. As illustrated in Fig. 1, the framework does not search the full detector space in a single pass. Instead, it fixes one module while optimizing the other, then switches roles and repeats this process iteratively. This design preserves search tractability for TinyML deployment while allowing the two modules to progressively co-adapt across iterations.
Within each stage shown in Fig. 1, the evolutionary search follows a standard population-based procedure, but is applied only to the currently active module. The population is initialized using two sources: high-performing candidates carried over from the module’s previous visit through Population Passthrough, and newly sampled candidates that maintain exploration. New architectures are then generated through crossover and mutation under the corresponding resource constraints (e.g., for the backbone and for the head). Each candidate is evaluated inside the weight-shared supernet using validation mAP, where the complementary module remains fixed. The best-performing candidates are retained to guide the next generation and to provide a strong starting point when the search returns to that module in a later cycle.
As depicted in Fig. 1, once the search for the current module is completed, ELASTIC switches to the other module and repeats the same procedure. By alternating in this back-and-forth manner, the framework incrementally refines both modules rather than optimizing them only once. This repeated revisiting is important because improving one module changes the context in which the other operates. Therefore, the iterative structure of ELASTIC enables the backbone and head to adapt to each other over time while remaining within deployment constraints.
@@ III-C Population Passthrough
| SVHN | Backbone-Only [4, 3, 20] | 75.53% | +0.18% | 10.8 hrs | 137.6 M | 0.73 M | 2.62 ms |
| Neck/Head-Only [7, 10, 24] | 75.35% | +0.00% | 4.7 hrs | 475.3 M | 1.01 M | 2.69 ms | |
@@
A key practical challenge in ELASTIC’s outer-loop cyclic module-wise optimization is maintaining optimization continuity when switching the active module (e.g., backbone neck/head backbone). Although the inner-loop evolutionary optimizer is standard, naive alternation effectively cold-starts each stage by discarding the module-specialized population, forcing repeated re-discovery of candidates compatible with the fixed context.
To make alternation workable, we introduce Population Passthrough, a lightweight memory mechanism that transfers elite architectural “knowledge” across module switches. For each module (backbone vs. neck/head), we maintain a passthrough buffer containing the top-performing candidates (ranked by validation mAP under the current resource constraints) from the most recent optimization of . When ELASTIC revisits , we initialize the population by mixing a fixed fraction of elites from (inheritance) with newly sampled/mutated candidates (diversity), balancing exploitation and exploration.
This mechanism is new in our setting because it targets alternating modular NAS: unlike standard elitism/aging schemes that assume a single, fixed search space and continuous run, Population Passthrough is applied at module boundaries to stabilize the outer-loop alternation and avoid repeated re-discovery.
@@ III-D Resource Allocation and Search Budget
In addition, ELASTIC is designed to support search interruption, allowing the iterative process to be paused at any stage and a valid, high-quality subnetwork to be extracted. This property is particularly beneficial in real-world TinyML development workflows, where compute resources are limited and iterative refinement may need to occur incrementally [11].
@@ IV-A Experimental Setup
We construct five OFA-based supernets tailored for diverse TinyML-relevant datasets: (i) SVHN, (ii) VGGFace2, (iii) a five-class subset of PascalVOC, (iv) the full PascalVOC dataset, and (v) TrashNet. These datasets are selected to cover a range of TinyML application domains, from lightweight digit recognition to face identification and general-purpose object detection, while avoiding the prohibitive memory and compute demands of large-scale datasets such as COCO.
The SVHN, VGGFace2, and PascalVOC subset supernets adopt a one-stage SSD-style architecture, while the full PascalVOC and TrashNet supernets adopt a MobileNet–YOLO style design that combines a lightweight MobileNet backbone with YOLO-style detection heads. For the SSD-style supernets, the search space varies (i) layer widths, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of layers allocated at each feature map. For both the MobileNet–YOLO supernets, the search space varies (i) expansion ratios, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of blocks allocated at each feature map.
@@
We evaluate deployment across three MCU platforms. The MAX78000/02 are accelerator-equipped microcontrollers with a dedicated CNN engine and separate on-chip memory for weights and activations. MAX78000 provides 442 kB weight memory and 524 kB data memory for the earliest streaming layers, while subsequent layers must fit within a single 32 kB memory instance. MAX78002 provides 2.4 MB weight memory and 1.3 MB data memory for the earliest streaming layers, while subsequent layers must fit within a single 80 kB memory instance. In contrast, the STM32F746 is a general-purpose microcontroller without a CNN accelerator, with 320 kB SRAM and 1 MB flash, where weights are stored in flash and activations are stored in SRAM. In this way, MAX78000/MAX78002 represent accelerator-equipped TinyML targets, while STM32F746 provides a non-accelerator MCU baseline for comparison.
Evolutionary search is performed with a population size of 100, mutation probability of 0.2, mutation ratio of 0.5, and a parent selection ratio of 0.25. Each search is budgeted for 55 generations on SVHN and VGGFace2, and 30 generations on the PascalVOC subset, full PascalVOC, and TrashNet datasets, respectively. Convergence is defined as the point at which the best candidate architecture’s mAP is within 1% of the final value achieved at the end of the search.
@@ IV-B Comparative Search Evaluation
To assess ELASTIC’s suitability for microcontrollers, we extend our experiments using the MCUNET [13] search space with the neck module included for fair comparison. Benchmarked on the full PascalVOC () dataset with similar parameter budgets (Table II), our ELASTIC-derived detector achieves 72.3% mAP, surpassing MCUNET by 20.9%, TinyissimoYOLO’s [15] best model by 16.3%, and MCUNetV2-M4/H7 [12] by 7.7% and 4.0%, respectively, while using 50–75% fewer MACs. Unlike MCUNetV2 [12], which requires external patch-based inference, ELASTIC supports native end-to-end inference, making it easier to deploy across diverse microcontrollers.
@@
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | N/A | 494.8 | N/A | 9.12 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | N/A | 490.6 | N/A | 19.0 |
| ai87-fpndetector [9] | TrashNet | MAX78002 | 2.18M | 62001 | 122.6 | 445.76 | 83.1 |
| ELASTIC (OURS) | TrashNet | MAX78002 | 1.32M | 17581 | 51.1 | 285.02 | 93.3 |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | N/A | 494.8 | N/A | 54.03 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | N/A | 490.6 | N/A | 73.9 |
@@
| ai85net-tinierssd-face [9] | VGGFace2 | STM32746 | 0.277M | N/A | 968.2 | N/A | 80.2 |
| ELASTIC (OURS) | VGGFace2 | STM32746 | 0.168M | N/A | 699.1 | N/A | 82.1 |
@@
| ai85net-tinierssd [9] | SVHN | STM32746 | 0.229M | N/A | 449.3 | N/A | 83.5 |
| ELASTIC (OURS) | SVHN | STM32746 | 0.235M | N/A | 523.7 | N/A | 87.9 |
IV-C Impact of Population Passthrough on Search Stability
Iterative NAS can suffer sharp drops at module boundaries because switching the active module effectively resets the evolutionary population. Without population memory, each alternation incurs a repeated “catch-up” phase to re-discover compatible high-quality candidates, harming stability and slowing convergence. We mitigate this with Population Passthrough, which carries a fixed fraction of elite architectures into the next stage to preserve optimization momentum.
Figure 3 ablates the passthrough ratio on PascalVOC. With 0% passthrough, mAP collapses after switches and saturates below 22%. Ratios of 50–80% stabilize the search and reach mAP in roughly half the iterations. 60% performs best, balancing exploitation (stability) and exploration (progress).